OpenStudy (anonymous):

1. Efficacy of a drug. Seventy-five (75) people with hypertension were given a new drug for the treatment of their condition. Blood pressure was lowered sufficiently to reduce the risk of serious outcomes in 20 of the 75 patients. (A) In what proportion of the sample did the drug work? (Calculate ).(2 points) (B) Calculate a 95% confidence interval for the proportion. Show all work, and interpret your findings. (Check to see if you can use a normal approximation before calculating the confidence interval.) (4 points) (C) How large a sample would be needed to reduce the margin of error to

OpenStudy (anonymous):

1)\[\hat p=\frac{ 20 }{ 75 }=0.27\] 2)For a proportion, confidence interval is defined as:\[\hat p-z_{\frac{ \alpha }{ 2 }}\sqrt{\frac{ \hat p(1-\hat p) }{ n }}\le p \le \hat p+z_{\frac{ \alpha }{ 2 }}\sqrt{\frac{ \hat p(1-\hat p) }{ n }}\]In our case:\[\hat p=0.27;z_{\frac{ 0.05}{ 2 }}=1.96;n=75\]